OpenStudy (anonymous):

ASAP PLZ HELP ASAP An expression is shown below: f(x) = -16x2 + 60x + 16 Part A: What are the x-intercepts of the graph of the f(x)? Show your work. Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph.

OpenStudy (danjs):

if y = f(x) = 0 then, it will cross the X axis there at points (x , 0)

OpenStudy (anonymous):

ok

OpenStudy (danjs):

y = 0 = -16x^2 + 60x + 16

OpenStudy (danjs):

maybe divide everything by 4 first

OpenStudy (anonymous):

ok

OpenStudy (danjs):

\[0 = 4x^2-15x+4\] i divided both sides by -4, can you factor that now?

OpenStudy (anonymous):

maybe

OpenStudy (danjs):

what you think... ?

OpenStudy (anonymous):

idk im bad with math :(

OpenStudy (anonymous):

4x^2−11?

OpenStudy (danjs):

You can expand that to check if it is right... 0 =(x - 4)* (4x +1)

OpenStudy (danjs):

So if either of those two quantities is Zero, then the equation is true... x - 4 = 0 or 4x+1 = 0 solve for the two x intercepts